OpenStudy (anonymous):

(last problem i need help with)Which of the following is a counterexample of the statement below? The product of two positive numbers is always greater than either number. a.2, 2 b.½, 2 c.3, 10 d.2, -1

OpenStudy (anonymous):

Well go through each option, and multiply the two values they give you. Then take the result of the multiplication, and see if either of the initial two values is greater than the product of the multiplication. I'll do the first one for you. \[\large \sf 2 \times 2 =4\] Both 2's are less than 4, so this proves the postulate right, which means it can't be the answer.

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

Well the postulate only applies to POSITIVE numbers, and since -1 is negative, we can't apply it.

OpenStudy (anonymous):

wait so is the 1/2 in the second option 0.5?

OpenStudy (anonymous):

In effect, yes. That is simply another way of representing it.

OpenStudy (anonymous):

if its not b then i have no idea

OpenStudy (anonymous):

Well yes, it is B. \[\large \sf \frac{1}{2} \times 2 =1\] \[\large \sf 2 > 1\] so the postulate is wrong, which is what our goal is.

OpenStudy (anonymous):

I understand this now! Thank you for helping me